# Scaffold-BPE: Enhancing Byte Pair Encoding with Simple and Effective Scaffold Token Removal

### v1

@@ Abstract
Byte Pair Encoding (BPE) serves as a foundation method for text tokenization in the Natural Language Processing (NLP) field. Despite its wide adoption, the original BPE algorithm harbors an inherent flaw: it inadvertently introduces a frequency imbalance for tokens in the text corpus. Since BPE iteratively merges the most frequent token pair in the text corpus while keeping all tokens that have been merged in the vocabulary, it unavoidably holds tokens that primarily represent subwords of complete words and appear infrequently on their own in the text corpus. We term such tokens as Scaffold Tokens. Due to their infrequent appearance in the text corpus, Scaffold Tokens pose a learning imbalance issue for language models. To address that issue, we propose Scaffold-BPE, which incorporates a dynamic scaffold token removal mechanism by parameter-free, computation-light, and easy-to-implement modifications to the original BPE. This novel approach ensures the exclusion of low-frequency Scaffold Tokens from the token representations for the given texts, thereby mitigating the issue of frequency imbalance and facilitating model training. On extensive experiments across language modeling tasks and machine translation tasks, Scaffold-BPE consistently outperforms the original BPE, well demonstrating its effectiveness and superiority.
[…]
284
@@ 1 Introduction
In recent years, large language models (LLM) [33, 8, 42] have made a substantial impact on Natural Language Processing (NLP). Those models, which are usually extremely huge in parameter scales [42, 34, 9], demand managing and processing vast vocabularies efficiently during both training and inference stages, and pose great challenges to the communities. To address that issue, Byte Pair Encoding (BPE) was applied to construct vocabularies. BPE was initially introduced as a data compression algorithm [14]. It iteratively merges the most frequent pairs of bytes or characters in a dataset until a desired vocabulary size is reached. The capability of BPE to break down words into more manageable subword units allows for more flexible and semantically complete representations of input data. Therefore, the BPE technique avoids the out-of-vocabulary problem [40], a prevalent issue where unknown words disrupt the functioning of NLP models. Consequently, BPE has been attracting much attention across the community [31, 49, 38]. In the NLP field, BPE is a cornerstone for numerous applications, like machine translation [40, 32, 49, 17], language understanding [27, 19], and even Large Language Model (LLM) training [33, 8, 5, 42, 50, 34, 52].
@@
Figure 1: Two types of tokens trained by the original BPE on the Pile dataset: one type, such as “age”, can appear frequently by itself, while the other type, such as “zona”, mostly appears as a subword of “Arizona” and thus has a low individual appearance frequency. The value on the horizontal axis represents the percentage of the frequency of the token relative to the total frequency of all tokens containing the specific token. For “age”, we only visualize the top tokens.
Since its inception, BPE has undergone various modifications to better suit the needs of complex natural language processing tasks. Existing works [45, 21, 12, 49, 17, 36, 39, 18, 32, 20, 38] have thoroughly investigated to improve BPE from multiple perspectives. A significant trend of research has focused on identifying the optimal vocabulary size that BPE should target [49, 17, 36, 39, 18, 32, 12]. These investigations reveal that a meticulously calibrated vocabulary size can reduce the computational load while improving the linguistic accuracy of the models. Besides, several works [32, 20, 38] have looked into the encoding process of BPE. By optimizing the encoding paths of tokens, these methods capture a wider range of linguistic phenomena and thus improve model performance across diverse datasets.
However, existing studies have overlooked a significant limitation inherent in the BPE method: the iterative merging process can lead to an imbalance in token frequencies by including low-frequency subwords as tokens in vocabulary. For example, as illustrated in Figure 1, in the commonly used Pile dataset [15] for training LLMs [5, 48, 41] tokenized by the original BPE, the token “zona” mostly appears as a subword within the token “Arizona” rather than as an independent, high-frequency token. Despite its lower standalone frequency, BPE includes “zona” in the final vocabulary because it is the “intermediate token” to derive the frequent token “Arizona”. We define such intermediate tokens that are crucial for constructing longer frequent tokens but do not appear frequently on their own as Scaffold Tokens. As we observe, a 32K vocabulary (size applied by LLaMA series [42, 43]), trained on the Pile dataset [15] with the original BPE [40], contains about 6.07% of scaffold tokens. As depicted in Figure 2, a natural frequency imbalance arises between these scaffold tokens and actual high-frequency tokens. Prior studies [26, 41] have highlighted that such disparities in token frequencies can result in imbalanced learning difficulties across different tokens. Scaffold tokens, due to their lower individual appearance frequencies, are notably harder to learn for models. To address that issue, we propose enhancements to the BPE algorithm aimed at mitigating the frequency imbalance and ensuring a more equitable learning process for all tokens.
@@
Specifically, we propose the simple and effective Scaffold-BPE with a dynamic scaffold tokens removal mechanism, which is parameter-free, computation-light, easy-to-implement, and widely effective. Generally, the proposed Scaffold-BPE expands the original BPE vocabulary. In the training stage, Scaffold-BPE dynamically marks tokens with lower individual appearance frequencies as scaffold tokens in each iteration. Note that the scaffold tokens are not actual tokens in the vocabulary and do not appear in the tokenized sequences after encoding. In the encoding stage, the Scaffold-BPE firstly utilizes all tokens in the expanded vocabulary to generate the token representations for the given texts, which is termed as a Scaffolding process. Then, the Scaffold-BPE ensures the absence of all scaffold tokens in the token representation by demolishing them into their shortest non-scaffold-tokens sequence, which is termed as a Demolishing process. Thanks to such modifications, Scaffold-BPE can remove scaffold tokens from the final token representations fed into models for more balanced token occurrences, thus leading to more sufficient learning and better performance of models.
We conduct extensive experiments on language modeling tasks and machine translation tasks. First, experiments on six widely used language modeling benchmarks including HellaSwag [51], OpenBookQA [29], PIQA [6], SIQA [37], StoryCloze [30], and Winogrande [35] demonstrate that Scaffold-BPE consistently outperforms the original BPE. Besides, experiments on WMT’14 English-German and WMT’14 English-France machine translation tasks [7] also indicate that Scaffold-BPE outperforms the original BPE.
@@
  • We observe that the iterative training process of BPE incorporates tokens with imbalanced frequency distributions into the vocabulary, which we term scaffold tokens.
@@
  • Extensive experiments demonstrate that Scaffold-BPE surpasses the original BPE on language modeling and machine translation tasks, proving its effectiveness and robustness in the NLP field.
2 Related Work
@@ 2.1 Byte Pair Encoding
In the early stages of neural model development, researchers primarily constructed models based on word-level vocabularies [11, 44, 53], which showed considerable success.
However, those models often struggled with the challenge of rare word handling due to the inherent limitations of word-level vocabulary size. In response, the academic community has explored numerous innovative strategies for vocabulary development, including methods based on bytes [46], characters [11, 25, 1], and subwords [40, 24].
Among those, Byte Pair Encoding (BPE) [40] stands out for its effective creation of subword vocabulary. Its design philosophy is notably straightforward. During the training process, the corpus is initially split into a sequence of the smallest unit tokens (i.e., character tokens [40] or byte tokens [46]). The algorithm iteratively finds the most frequent token pairs in the sequence, merges them into a new token, and adds it to the vocabulary until it reaches a predetermined size. The vocabulary is then utilized during the encoding phase to represent any text. It reduces token sparsity and enhances feature identification in related words sharing an identical subword, without losing rare words. Recent advancements like BPE-dropout [32] and optimal vocabulary size search [49, 17, 39, 36] continue to enrich BPE development in neural models.
However, previous works did not take into account a fundamental flaw of BPE: during the training process of BPE, existing tokens are merged into new tokens, resulting in a decrease of their individual frequencies that are not covered by the new tokens. That leads to the inclusion of some lower-frequency tokens in the vocabulary, hindering the inclusion of other actual high-frequency tokens, thus resulting in an imbalance of token frequencies and wastage of the vocabulary. To address that issue, this paper introduces Scaffold-BPE, which has a dynamic scaffold token removal mechanism that ensures the tokens fed into models are actual high-frequency tokens.
2.2 Language Models
Language models are designed to predict the probability distribution of a token sequence. Following GPT-3 [8], which features 175 billion parameters and demonstrates versatility across a wide range of applications, there has been a significant push towards developing large generative language models like Gopher [34], PaLM [9], GaLM [13], OPT [52], and LLaMA [42]. Such a surge in development has greatly advanced the fields of natural language understanding and generation.
However, as a mainstream tokenizer used by language models [33, 8, 5, 42, 50, 34, 52], BPE has inherent flaws of imbalanced token frequencies, which can impact the training of language models. First, the low-frequency tokens are updated infrequently during the training process, leading to poorer performance of those tokens. Second, the presence of low-frequency tokens in the vocabulary prevents other actual high-frequency tokens from being included. Together, the two points above further impact the performance of language models on downstream tasks. This paper demonstrates that by using our Scaffold-BPE algorithm as the tokenizer, language models can achieve a consistent improvement on downstream tasks.
@@ 3 Methodology
To enhance the performance of the original BPE, we propose Scaffold-BPE to remove the scaffold tokens introduced by the original BPE. Our Scaffold-BPE is simple and straightforward. In the training process, the Scaffold-BPE dynamically marks scaffold tokens in the vocabulary at each iteration, and finally yields an expanded vocabulary consisting of both normal tokens with the amount equaling the predetermined vocabulary size and several scaffold tokens. In the encoding process, apart from using the normal tokens, Scaffold-BPE temporarily uses previously marked scaffold tokens as intermediate tokens to merge into longer normal tokens.
@@ 3.1 Training Process
Algorithm 1 Scaffold-BPE Training Algorithm
Text Corpus $C$, Desired Vocabulary Size $N$
Initialize an expanded vocabulary $E$, consisting of a normal-token vocabulary $V$ and a scaffold-token vocabulary $S$
Split $C$ into a list of characters/bytes (denoted as $L$)
Initialize a priority queue $Q$(token pair, $f(\cdot)\downarrow$) within $L$
while $|V|$ < $N$ do
     $(a,b)$leftarrow pop $Q_{head}$
     Merge pair $(a,b)$ into a new token $t$
     if $t$ in $S$ then $\triangleright$ $t$ may be a previously marked scaffold token
         Scaffold$(t)\leftarrow$ False
         continue
     end if
     Add $t$ to $E$ as a normal token
     Replace all of $(a,b)$ in $L$ with $t$
     Update $Q$
     /******* Scaffold-BPE Modification Begins *******/
     for each $p$ in {$a$, $b$do
         $f(p)\leftarrow f(p)-f((a,b))$
         if $p$ in $V$ and $f(p)$ < $f(Q_{head})$ then
              Scaffold$(t)\leftarrow$ True
              Push $p$ back to $Q$
         end if
     end for
     /******** Scaffold-BPE Modification Ends ********/
end while
return $E$ (with $V$ and $S$ both included)
@@
BPE then merges (i.e., concatenates) them into a new token $t$, and includes $t$ in $V$. Then BPE updates $L$ via replacing all $(a,b)$ with $t$, and restarts the process again. The iterative process of identifying the most frequent token pair $(a,b)$ can be accelerated using a priority queue $Q$. At the beginning of the training process, all token pairs in $L$ are pushed into $Q$ with a descending order of frequency. And after the token pair $(a,b)$ is merged into $t$ in each iteration, BPE updates the frequency and rank of token pairs related to all indexed occurrences of $(a,b)$. For instance, given $(a,b)$ in a context of $\ldots,u,a,b,v,\ldots$ in $L$, when $(a,b)$ is replaced with $t$, the frequency of $(u,a)$ or $(b,v)$ would decrease by 1, and meanwhile that of $(u,t)$ or $(t,v)$ would increase by 1. With the occurrences of all token pairs being indexed, there is no need to scan $L$ again and re-count the frequencies of all candidate token pairs for a new iteration. After updating the adjacent token pairs related to $(a,b)$ (i.e, $t$), the frequencies of token pairs like $(u,a)$ or $(b,v)$ would be updated in $Q$, and meanwhile the new candidate token pairs $(u,t)$ and $(t,v)$ would also be pushed into $Q$ with their corresponding frequencies.
The Scaffold-BPE expands the vocabulary $V$ to an expanded vocabulary $E$, and assigns an attribute (denoted as $\text{Scaffold}(\cdot)$) to each token in the vocabulary indicating whether it is a scaffold token or not. Thus, the expanded vocabulary $E$ comprises two types of tokens. We denote all the non-scaffold-tokens by $V$, which, as with the original BPE, are the tokens actually used in model training.
@@
Additionally, we denote all the scaffold tokens by $S$, which are not fed into the model, nor do they appear in any token representations after encoding.
@@
They only serve as intermediate tokens to aid in the training and encoding processes of Scaffold-BPE. Therefore, when calculating the size of the vocabulary, the count of scaffold tokens is not included; only the number of tokens in $V$ is considered.
Initially, a token pair is merged and added to $E$ due to its high frequency. Similarly, Scaffold-BPE marks a token as a scaffold token when its frequency decreases. Throughout the entire training process of BPE, $f(a)$ and $f(b)$ only decrease when the token pair $(a,b)$ is merged into a new token $t$. Therefore, as presented in Algorithm 1, Scaffold-BPE introduces an additional step at the end of each iteration, utilizing the reduced $f(a)$ and $f(b)$ to evaluate whether $a$ and $b$ remain high-frequency. If they are no longer considered high-frequency, they would be marked as scaffold tokens. Naturally, the token pair at the head of the priority queue $Q$ (denoted as $Q_{head}$) is the next candidate to be added to the vocabulary. Then $f(Q_{head})$ is a natural frequency delimiter between in-vocabulary and out-vocabulary tokens. Therefore, if $f(a)$ (or $f(b)$) $<f(Q_{head})$, $a$ (or $b$) is marked as a scaffold token, which means it is not included by $V$. Notably, such an additional step leverages the inherent mechanism of BPE without introducing any additional hyper-parameters, maintaining the simplicity and clarity of BPE. Moreover, $f(Q_{head})$ is dynamically adjusted in each iteration, ensuring that Scaffold-BPE can adaptively identify scaffold tokens at any iteration step. Furthermore, scaffold tokens are not permanently marked. They are pushed back into $Q$, reserving the possibility of being ranked top at the priority queue and re-integrated into $V$ in a future iteration.

### v2

# Scaffold-BPE: Enhancing Byte Pair Encoding for Large Language Models with Simple and Effective Scaffold Token Removal

@@ Abstract
Byte Pair Encoding (BPE) serves as a foundation method for text tokenization in the Natural Language Processing (NLP) field. Despite its wide adoption, the original BPE algorithm harbors an inherent flaw: it inadvertently introduces a frequency imbalance for tokens in the text corpus. Since BPE iteratively merges the most frequent token pair in the text corpus to generate a new token and keeps all generated tokens in the vocabulary, it unavoidably holds tokens that primarily act as components of a longer token and appear infrequently on their own. We term such tokens as Scaffold Tokens. Due to their infrequent occurrences in the text corpus, Scaffold Tokens pose a learning imbalance issue. To address that issue, we propose Scaffold-BPE, which incorporates a dynamic scaffold token removal mechanism by parameter-free, computation-light, and easy-to-implement modifications to the original BPE method. This novel approach ensures the exclusion of low-frequency Scaffold Tokens from the token representations for given texts, thereby mitigating the issue of frequency imbalance and facilitating model training. On extensive experiments across language modeling and even machine translation, Scaffold-BPE consistently outperforms the original BPE, well demonstrating its effectiveness.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have become a burgeoning paradigm in handling a broad array of Natural Language Processing (NLP) tasks. The tokenization process in most modern LLMs (Radford et al. 2019; Brown et al. 2020; Rae et al. 2021; Zhang et al. 2022; Biderman et al. 2023; Touvron et al. 2023a; Yang et al. 2023; Achiam et al. 2023; Dubey et al. 2024) employs Byte Pair Encoding (BPE) (Sennrich et al. 2015), a method that was originally designed for data compression (Gage 1994). BPE consists of two main stages. In the training stage, BPE iteratively merges the most frequent pairs of bytes or characters in a dataset into a new token and adds it to the vocabulary until a desired vocabulary size is reached. And in the encoding stage, the vocabulary is utilized to represent any text. The adoption of BPE in LLMs is driven by its capability to decompose words into smaller, manageable subword units, thus avoiding out-of-vocabulary words, facilitating flexible and semantically complete representations of input data. Actually, BPE has also been widely used in traditional NLP tasks, like machine translation (Provilkov et al. 2019; Xu et al. 2020), sentence classification (Liu et al. 2019; He et al. 2020b) and summarization (Wu et al. 2021; Xu et al. 2022).
@@
Figure 1: Two types of tokens in original BPE vocabulary on the Pile dataset: one type, such as “ing”, appears frequently by itself, while the other type, such as “zona”, mostly appears as a component of “Arizona” and thus has a low individual occurrence frequency. The value on the horizontal axis represents the percentage of the frequency of the token relative to the total frequency of all tokens containing “ing”/“zona”. For “ing”, we only visualize the top tokens.
Since its inception, BPE has undergone various modifications to better suit the needs of complex NLP tasks, including identifying the optimal vocabulary size for various tasks (Xu et al. 2020; Gutierrez-Vasques et al. 2021), optimizing the encoding paths of tokens to achieve subword regularization (Provilkov et al. 2019; He et al. 2020a), etc.
@@
However, existing studies have overlooked an inherent limitation in the BPE method: the iterative merging process can lead to an imbalance in token frequencies by including low-frequency tokens in vocabulary. For example, as illustrated in Figure 1, in the commonly used Pile dataset (Gao et al. 2020) that is tokenized by the original BPE method of 32K vocabulary size (as LLaMA series (Touvron et al. 2023a, b)), the token “zona” mostly appears as a component of the token “Arizona” rather than as an independent, high-frequency token. Despite its lower standalone frequency, BPE includes “zona” in the final vocabulary because it is the “intermediate token” to derive the frequent token “Arizona”. We define such intermediate tokens that are crucial for constructing longer frequent tokens but do not appear frequently on their own as Scaffold Tokens. Note that not all subwords are simply scaffold tokens. For example, “ing” is not identified as a scaffold token, as there are many words containing “ing” but are not tokens in the vocabulary. For example, “connecting” is represented as “connect”+“ing” (2 tokens). Such words help to keep “ing” a frequent token. Therefore, “ing” is not a scaffold token. According to our proposed Scaffold-BPE method, the 32K vocabulary contains about 6.07% of scaffold tokens.
As depicted in Figure 2, a natural frequency imbalance arises between these scaffold tokens and actual high-frequency tokens. Prior studies (Lin et al. 2017; Su et al. 2023) have highlighted that such disparities in token frequencies can result in imbalanced learning difficulties across different tokens. Scaffold tokens, due to their lower individual occurrence frequencies, are notably harder to learn for models.
To address that issue, we propose enhancements to the BPE algorithm, aiming to mitigate the frequency imbalance and ensure a more equitable learning process for all tokens. Specifically, we propose the simple and effective Scaffold-BPE with a dynamic scaffold token removal mechanism, which is parameter-free, computation-light, easy-to-implement, and widely effective. Generally, the proposed Scaffold-BPE maintains an expanded vocabulary compared with the original BPE, which consists of both normal tokens and scaffold tokens. Note that the scaffold tokens are not actual tokens in the vocabulary and do not appear in the tokenized sequences after encoding. In the training stage, Scaffold-BPE dynamically marks tokens with lower individual occurrence frequencies as scaffold tokens in each iteration. In the encoding stage, the Scaffold-BPE firstly utilizes all tokens in the expanded vocabulary to generate the token representations for the given texts, which is termed as a Scaffolding process. Then, the Scaffold-BPE ensures the absence of all scaffold tokens in the token representation by demolishing them into their shortest non-scaffold-token sequence, which is termed as a Demolishing process. Thanks to such modifications, Scaffold-BPE can remove scaffold tokens from the final token representations fed into models, thus enjoying more balanced token occurrences, leading to more sufficient learning and better performance of models.
We conduct extensive experiments on language modeling tasks. Results on 9 widely used language modeling benchmarks demonstrate that Scaffold-BPE consistently outperforms the original BPE. Besides, even when extended to machine translation tasks, Scaffold-BPE proves highly effective. Furthermore, we show that Scaffold-BPE is orthogonal to existing modifications on BPE, like BPE-Dropout (Provilkov et al. 2019) and can be combined with them to achieve further improvements.
@@
  • We observe that the iterative training process of BPE incorporates low-frequency tokens into the vocabulary, which we term scaffold tokens.
@@
  • Extensive experiments demonstrate that Scaffold-BPE outperforms the original BPE on language modeling and also machine translation tasks, proving its effectiveness and robustness in the NLP field.
2 Related Works
Recently, LLMs have become a popular paradigm for solving NLP tasks, with BPE serving as the mainstream tokenizer to split a text into a sequence of tokens (e.g., subwords/words/phrases). Thus, enhancing BPE could boost the performance of LLMs and have positive implications for various applications.
2.1 Language Models
Language models are designed to maximize the likelihood of a token sequence. Following GPT-3 (Brown et al. 2020), which features 175 billion parameters and demonstrates versatility across a wide range of applications, there has been a significant push towards developing large generative language models like Gopher (Rae et al. 2021), PaLM (Chowdhery et al. 2023), GaLM (Du et al. 2022), OPT (Zhang et al. 2022), and LLaMA (Touvron et al. 2023a). Such a surge in development has greatly advanced the fields of natural language understanding and generation.
@@ 2.2 Byte Pair Encoding
Early neural models had difficulty managing rare words due to limited vocabulary sizes. BPE (Sennrich et al. 2015) effectively addresses that by generating a subword vocabulary. Initially, a corpus is split into characters or bytes, which act as initial tokens. The algorithm iteratively finds the most frequent token pair in the sequence, merges them into a new token, and adds it to the vocabulary until it reaches a predetermined size. The vocabulary is then utilized during the encoding phase to represent any text. Recent advancements like BPE-dropout (Provilkov et al. 2019) and optimal vocabulary size search (Xu et al. 2020; Gowda et al. 2020; Salesky et al. 2020) continue to enrich BPE developments.
However, previous works did not take into account a fundamental flaw of BPE: the iterative training process of BPE incorporates low-frequency tokens into the vocabulary, hindering the inclusion of other actual high-frequency tokens, thus resulting in an imbalance of token frequencies and wastage of the vocabulary. To address that issue, here we propose Scaffold-BPE, which has a dynamic scaffold token removal mechanism that ensures the tokens fed into models are always actual high-frequency tokens.
@@ 3 Methodology
Algorithm 1 Scaffold-BPE Training Algorithm
0:  Text Corpus $C$, Desired Vocabulary Size $N$
1:  Initialize an expanded vocabulary $E$, consisting of a normal-token vocabulary $V$ and a scaffold-token vocabulary $S$
2:  Split $C$ into a list of characters/bytes (denoted as $L$)
3:  Initialize a priority queue $Q$ storing token pairs within $L$, arranged in reverse order of frequency
4:  while $|V|<N$ do
5:     $(a,b)$leftarrow pop $Q_{head}$
6:     Merge pair $(a,b)$ into a new token $t$
7:     if $t$ in $S$ then
8:        // $t$ may be a previously marked scaffold token
9:        Scaffold$(t)\leftarrow$ False
10:        continue
11:     end if
12:     Add $t$ to $E$ as a normal token
13:     Replace all of $(a,b)$ in $L$ with $t$
14:     Update $Q$
15:     /****** Scaffold-BPE Modification Begins ******/
16:     for each $t^{\prime}$ in {$a$, $b$do
17:        $f(t^{\prime})\leftarrow f(t^{\prime})-f(t)$
18:        if $t^{\prime}$ in $V$ and $f(t^{\prime})<f(Q_{head})$ then
19:           Scaffold$(t^{\prime})\leftarrow$ True
20:           Push $t^{\prime}$ back to $Q$
21:        end if
22:     end for
23:     /****** Scaffold-BPE Modification Ends ******/
24:  end while
25:  return  $E$ (with $V$ and $S$ both included)
Algorithm 2 Scaffold-BPE Encoding Algorithm
0:  A Text $T$, Expanded Vocabulary $E$
1:  Split $T$ into a character/byte token representation (denoted as $R$)
2:  while True do
3:     /************ Scaffolding Begins ************/
4:     Identify all possible merges $M$ using $E$, ignoring token types
5:     /************* Scaffolding Ends *************/
6:     if $M$ is empty then
7:        break
8:     end if
9:     Select $m$ which is ranked before the others in $E$ from $M$
10:     Apply $m$ to $R$
11:  end while
12:  /************ Demolishing Begins ************/
13:  Demolish all scaffold tokens in $R$ into its shortest non-scaffold child token sequence
14:  /************* Demolishing Ends *************/
15:  return  $R$
To enhance the original BPE, we propose Scaffold-BPE to remove the scaffold tokens introduced by the original BPE. Our Scaffold-BPE is simple yet effective. In the training process, the Scaffold-BPE dynamically marks scaffold tokens in the vocabulary at each iteration, and finally yields an expanded vocabulary consisting of both normal tokens with the amount equaling the predetermined vocabulary size and several scaffold tokens. In the encoding process, apart from using the normal tokens, Scaffold-BPE temporarily uses scaffold tokens as intermediate tokens to merge into longer normal tokens.
@@ 3.1 Training Process
BPE then merges (i.e., concatenates) them into a new token $t$, and includes $t$ in $V$. Then BPE updates $L$ via replacing all $(a,b)$ with $t$, and restarts the process again.
The iterative process of identifying the most frequent token pair $(a,b)$ can be accelerated using a priority queue $Q$. At the beginning of the training process, all token pairs in $L$ are pushed into $Q$ with a descending order of frequency. And after the token pair $(a,b)$ is merged into $t$ in each iteration, BPE updates the frequencies and the ranks of token pairs related to all indexed occurrences of $(a,b)$. For instance, given $(a,b)$ in a context of “$\ldots,u,a,b,v,\ldots$” in $L$, when $(a,b)$ is replaced with $t$, the frequency of $(u,a)$ or $(b,v)$ would decrease by 1, and meanwhile that of $(u,t)$ or $(t,v)$ would increase by 1. With the occurrences of all token pairs being indexed, there is no need to scan $L$ again and re-count the frequencies of all candidate token pairs for a new iteration. After updating the adjacent token pairs related to $(a,b)$ (i.e, $t$), the frequencies of token pairs like $(u,a)$ or $(b,v)$ would be updated in $Q$, and meanwhile the new candidate token pairs $(u,t)$ and $(t,v)$ would also be pushed into $Q$ with their frequencies.
The Scaffold-BPE expands the vocabulary $V$ to an expanded vocabulary $E$, and assigns an attribute (denoted as $\text{Scaffold}(\cdot)$) to each token in the vocabulary indicating whether it is a scaffold token or not. Thus, the expanded vocabulary $E$ comprises two types of tokens, i.e., normal ones and scaffold ones. We denote all the non-scaffold tokens by $V$, which, as with the original BPE, are the tokens actually used in representing texts for NLP model training:
@@
Additionally, we denote all the scaffold tokens by $S$, which are not fed into the model, nor do they appear in any token representations after encoding:
@@
They only serve as intermediate tokens to aid in the training and encoding processes of Scaffold-BPE. Therefore, when calculating vocabulary size, the count of scaffold tokens is not included; only the number of tokens in $V$ is considered.
BoolQ HellaSwag OpenBookQA PIQA SIQA StoryCloze Winogrande
468M Original BPE 58.64 40.78 30.50 66.57 43.40 62.77 53.00
Scaffold-BPE 60.52 41.68 32.20 68.69 44.09 63.04 54.22
1.2B Original BPE 60.86 47.25 31.70 68.55 44.09 65.61 55.52
Scaffold-BPE 62.26 48.07 32.90 69.86 45.34 67.02 56.00
6.7B Original BPE 62.87 60.57 35.10 73.69 46.98 71.43 60.97
Scaffold-BPE 64.95 61.19 38.00 74.54 47.49 72.26 61.76
Table 1: At varying model scales, the average accuracy on 0/5-shot common sense reasoning benchmarks ($p$-value $<0.01$).
Initially, a token pair is merged and added to $E$ due to its high frequency. Similarly, Scaffold-BPE marks a token as a scaffold token when its frequency decreases too much. Throughout the entire training process of BPE, $f(a)$ and $f(b)$ only decrease when the token pair $(a,b)$ is merged into a new token $t$. Therefore, as presented in Algorithm 1, Scaffold-BPE introduces an additional step at the end of each iteration, utilizing the decreased $f(a)$ and $f(b)$ to evaluate whether $a$ and $b$ remain high-frequency. If they are no longer considered high-frequency, they would be marked as scaffold tokens. Naturally, the token pair at the head of the priority queue $Q$ (denoted as $Q_{head}$) is the next candidate to be added to the vocabulary. Then $f(Q_{head})$ is a natural frequency delimiter between in-vocabulary and out-of-vocabulary tokens. Therefore, if $f(a)$ (or $f(b)$) $<f(Q_{head})$, $a$ (or $b$) is marked as a scaffold token, which means it is not included in $V$:
$\text{Scaffold}(a)=\begin{cases}\text{True},&\text{if }f(a)<f(Q_{head})\\ \text{False},&\text{otherwise}\end{cases}$ (4)
Notably, such an additional step leverages the inherent mechanism of BPE without introducing any additional hyper-parameters, maintaining the simplicity and clarity of BPE. Moreover, $f(Q_{head})$ is dynamically adjusted in each iteration, ensuring that Scaffold-BPE can adaptively identify scaffold tokens at any iteration step. Furthermore, scaffold tokens are not permanently marked. They are pushed back into $Q$, reserving the possibility of being ranked top at the priority queue and re-integrated into $V$ in a future iteration.